# Pro-Pose: Unpaired Full-Body Portrait Synthesis via Canonical UV Maps

### v1

@@ 3.4 Training Data and Conditioning Strategy
We enable joint training via a unified condition vector $\mathbf{c}=\{\mathbf{T}_{\text{in}},\mathbf{p}_{\text{target}},\mathbf{I}^{FC}_{\text{in}}\}$, comprising the input texture, target pose, and an optional identity-anchoring face crop extracted via MediaPipe [27]. An overview of our unified training framework is presented in Figure 3. We instantiate $\mathbf{c}$ based on the data source:
@@ 3.5 Test-Time Personalization via Few-Shot Adaptation
We constrain this personalization to the face by applying the loss exclusively to the facial region using a binary mask $\mathbf{M}_{j}$, which represents the target’s latent-space face mask:
@@
where $\mathbf{c}_{\text{paired}}=\{\mathbf{T}_{i},\mathbf{p}_{j},\mathbf{I}^{FC}_{i}\}$ is the conditions extracted from a sample pair $(i,j)$ drawn from $\mathcal{S}$ (such that $i\neq j$), and $v_{\text{target}}=\mathbf{x}_{j,1}-\mathbf{x}_{j,0}$ is the target transport guidance.
@@ 3.6 Implementation Details
Figure 5: Base Clothing (BC) Standardization. We apply different preprocessing strategies based on the data. For DeepFashion [25] and Commerce images [56, 55], we generate the base garment while preserving pose and identity via pixel-aligned editing (Prompt V1). For FFHQ [16], we use generative outpainting to expand limited face crops into full-body samples (Prompt V2).
Basic Clothing (BC) Dataset. We construct a standardized training dataset by processing paired DeepFashion [25] data, unpaired FFHQ [16] faces, and commercial images [56, 55] using Gemini 2.5 Flash Image [7]. As shown in Figure 5, we employ distinct strategies: pixel-aligned editing for full-body inputs to standardize garments to a black tank top and shorts, and generative outpainting to expand face crops into full-body portraits. The specific text prompts are detailed in the Appendix.
DeepFashion (In-Domain)
[…]
CFLD [26] 17.645 7.149 0.7478 0.1819 0.48 0.3180 0.9731 4.15
MCLD [24] 18.211 7.079 0.7557 0.1757 0.49 0.3440 0.9654 4.29
LEFFA [53] 14.020 4.23 0.7550 0.1190 0.44 0.5794 0.9409 4.41
OnePoseTrans [6] 13.570 8.740 0.6053 0.3066 0.46 0.5750 0.9476 4.32
UniHuman [20] 14.050 6.250 0.7960 0.1560 0.46 0.5810 0.9434 4.03
Unpaired Only 15.560 7.780 0.7013 0.2455 0.47 0.3400 0.9189 4.23
Paired Only 19.350 4.23 0.8094 0.0760 0.48 0.6103 0.9757 7.25
Ours (Unpaired + paired) 19.330 4.24 0.8095 0.0780 0.48 0.5919 0.9712 7.23
WPose (Out-of-Domain)
[…]
CFLD [26] 15.43 96.07 0.744 0.208 0.31 0.0885 0.6412 1.94
MCLD [24] 15.64 94.23 0.759 0.201 0.35 0.0995 0.6478 1.96
LEFFA [53] 16.71 67.85 0.776 0.193 0.32 0.0914 0.5725 2.01
OnePoseTrans [6] 17.23 27.43 0.818 0.151 0.33 0.1735 0.7205 4.44
UniHuman [20] 17.64 27.748 0.807 0.161 0.34 0.1121 0.7207 2.89
Unpaired Only 16.06 6.98 0.7565 0.2179 0.37 0.3800 0.6748 4.26
Paired Only 18.23 6.73 0.8143 0.1615 0.34 0.4954 0.6915 7.23
Ours (Unpaired + paired) 19.87 6.04 0.849 0.127 0.38 0.5563 0.7383 7.49
Table 1: Quantitative evaluation and ablation study. We report results on DeepFashion (top) and WPose (bottom). The upper section of each block compares our method against state-of-the-art baselines, while the lower section details our ablation study (Unpaired Only, Paired Only, and our full hybrid model). Red indicates the best performance and Blue indicates the second best across all methods.
Training Setup. Our combined dataset totals approximately 500k samples: 470K single unpaired images [16, 56, 55], and 30K paired images from DeepFashion dataset[25] (We restrict training to a subset of DeepFashion data to prevent overfitting to its limited number of unique identities ($\approx 100$), while ensuring all identities are still represented.)
Optimization. We use Flux.1 [dev] [18] as our backbone, training rank-128 LoRA adapters using AdamW with a learning rate of $10^{-4}$. Training proceeds for 100K iterations on 128 TPUv5 chips with a global batch size of 128.
@@ 4 Experiments
Base 18.59 6.88 0.823 0.4837 0.689 7.15
Finetuned 19.34 6.76 0.835 0.5722 0.705 7.17
Table 2: Ablation study (Finetuning). Finetuning can lead to better face and identity preservation compared to our base model.
4.1 Datasets and Metrics for Evaluation
Following prior work [24, 26, 53], we use the DeepFashion In-Shop Clothes Retrieval test split [25] (8570 image pairs). To assess generalization to in-the-wild scenarios, we additionally evaluate on WPose dataset [20] (2305 image pairs).
Since our method canonicalizes clothing, direct comparison against ground-truth images (which include diverse garments and backgrounds) is not meaningful. For image similarity (PSNR, SSIM [45]) and perceptual metrics (FID [10], LPIPS [51]), we therefore compare our outputs against a Basic Clothing version. We further report these metrics on the original image only face regions in the appendix.
Pose, Identity, Semantic and Perceptual Metrics. To measure how well the generated avatar matches the target pose, we compute Object Keypoint Similarity (OKS) [22] between the predicted and ground-truth keypoints. Identity fidelity is evaluated within the facial region. We compute Face Similarity (FaceSim) using cosine similarity of ArcFace [4] identity embeddings. We report DINOv2 similarity [35] as a measure of semantic and structural alignment. For reference-free perceptual quality, we report HPSv3 [30].
Figure 7: Finetuning Qualitative results. Although the input image suffered from motion blur, leading to blurry base model results, finetuning successfully restored crisp, detailed face features, demonstrating the model’s ability to learn robust features from multiple images rather than relying solely on one input.
4.2 Quantitative and Qualitative Comparison
We compare our method against the recent diffusion-based pose-conditioned avatar generation models, including MCLD [24], CFLD [26], LEFFA [53], OnePoseTrans [6], and UniHuman [20].
As detailed in Table 1, our method achieves SOTA performance on both benchmarks. On the in-domain DeepFashion set [25], our model sets a new SOTA, significantly outperforming all baselines across image fidelity (PSNR, SSIM, LPIPS), identity (FaceSim), and perceptual quality (HPSv3). For the in-the-wild WPose dataset [20], where we use foreground-masked metrics (M-PSNR, M-SSIM, M-LPIPS), our full model (Unpaired + Paired) conclusively dominates all SOTA baselines. This highlights the critical role of our unpaired data strategy in achieving robustness for challenging, in-the-wild data.
Figure 9 provides a qualitative comparison with SOTA methods. As demonstrated, our method generates high-fidelity avatars that accurately preserve the input person’s identity and facial features. For instance, the eighth row shows robust preservation of face and beard shape. Furthermore, our approach more robustly follows the target pose than any other method (e.g., sixth row) and maintains body shape fidelity relative to the original image (e.g., ninth row).
4.3 Ablation Study
@@
As shown in Table 1, while the Paired Only model performs well on in-domain data (DeepFashion [25]), it suffers on out-of-domain the WPose set [20], confirming its poor generalization. Figure 6 demonstrates that training solely on the limited paired dataset leads to overfitting and identity leaking from the training data. Alternatively, the model trained exclusively on unpaired data fails when subjected to significant pose changes. Our full method (Unpaired + Paired) synergizes these sources, effectively utilizing the larger dataset to learn robust identity priors while maintaining stable pose guidance.
@@
Figure 8: Limitations of the base model. Left: Extreme pose changes (e.g., back-to-front) result in minimal UV texture overlap, forcing the model to hallucinate unseen regions. While general attributes like beard shape or hair color may be inferred from partial cues, precise facial structure can lose fidelity. Right: Because our explicit UV conditioning is restricted to the face, body-specific skin details such as tattoos are not preserved.
Figure 9: Qualitative Comparison. The first two rows display results from the DeepFashion dataset [25], and the subsequent rows utilize real images from the WPose dataset [20]. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [24], Leffa [53], UniHuman [20], OnePoseTrans [6]) compared to Ours and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, particularly the facial characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result.
@@ 5 Limitations and Future Work
While our framework significantly advances single-image avatar reposing, some limitations remain. First, the base model can struggle with extreme pose changes (Figure 8, left), where minimal texture overlap forces identity-reducing hallucination. Although multi-image fine-tuning mitigates this issue, the base model may still fail in challenging cases. Second, our UV conditioning is restricted to the face; thus, distinctive body features like tattoos (Figure 8, right) or birthmarks cannot be represented on generic skin regions. Future work will address these limitations by extending UV conditioning to full-body skin texture and applying the skin- preservation during finetuning stage.
@@ 6 Conclusion
In this paper, we presented a novel framework for generating photorealistic, reposed human avatars from unconstrained single images. By introducing a self-supervised Donor-based UV Reposing strategy, we effectively decoupled pose from texture, enabling our Flow Matching generator to learn robust identity preservation from massive unpaired image collections. We further showed that this strong base model can be rapidly specialized via multi-image fine-tuning to handle more extreme reposing scenarios. By standardizing data into a base clothing and formulating generation in canonical UV space, our approach provides a scalable foundation for creating high-fidelity, controllable digital avatars from everyday photography.
[…]
@@
Figure 10: Pro-Pose In-the-wild. We apply our method to real-world images characterized by varying poses and environments. The columns display (from left to right): (1) the original input image; (2) the target pose represented by a SMPL-X mesh (reposed using ground truth parameters); (3) the output generated by Pro-Pose with a corresponding face crop for detailed visual comparison. Our method consistently preserves facial details and structure compared to the ground truth.
7 Base Clothing (BC) Dataset.
We utilize paired data from DeepFashion [25], unpaired single images from FFHQ [16], and commercial data derived from e-commerce websites [56, 55]. To ensure visual uniformity across these diverse sources, we generate the Base Clothing (BC) dataset using an identity-preserving I2I model (Gemini 2.5 Flash Image [7]), as shown in Figure 5 (main paper). Depending on the nature of the input data, we employ either Prompt V1 or Prompt V2:
@@
Precisely edit IMAGE_0 to change the garment worn by the person in the photo to a black sleeveless tank top and black shorts. Do not change the person in the image or the framing. Remove or replace any accessories such as bags, hats, jewelry and sunglasses in IMAGE_0. Do not alter IMAGE_0 in any other way. Keep it pixel-aligned.
@@
8 Canonical UV-Space Formulation

### v2

@@ 3.4 Training Data and Conditioning Strategy
We enable joint training via a unified condition vector $\mathbf{c}=\{\mathbf{T}_{\text{in}},\mathbf{p}_{\text{target}},\mathbf{I}^{FC}_{\text{in}}\}$, comprising the input texture, target pose, and an optional identity-anchoring face crop extracted via MediaPipe [MediaPipe]. An overview of our unified training framework is presented in Figure 3. We instantiate $\mathbf{c}$ based on the data source:
@@ 3.4.1 Paired Supervision
Figure 3: Overview of our Avatar Generation Framework. Our approach leverages single-view datasets by operating in a canonical UV space, extracting UV texture and pose [Pavlakos2019SMPLX]. Left (Paired Supervision): When ground-truth pose pairs are available, we condition the Flow Matching model on the partial UV texture, target pose, and face crop. Right (Single-View Self-Supervision): To prevent "pose leakage" from occlusion boundaries when training on single images, we introduce a Donor-based UV Reposing module (Sec. 3.2.1). This synthetically re-poses the input texture using a random donor visibility mask, forcing the model to learn robust identity representations. Furthermore, we drop out the face crop condition in this branch to prevent trivial reconstruction via pixel-perfect information leakage.
@@ 3.5 Test-Time Personalization via Few-Shot Adaptation
We apply the personalization loss to the full visible foreground using a binary mask $\mathbf{M}_{j}$, which represents the target’s latent-space skin segmentation mask:
[…]
@@
where $\mathbf{c}_{\text{paired}}=\{\mathbf{T}_{i},\mathbf{p}_{j},\mathbf{I}^{FC}_{i}\}$ are the conditions extracted from a sample pair $(i,j)$ drawn from $\mathcal{S}$ (such that $i\neq j$), and $v_{\text{target}}=\mathbf{x}_{j,1}-\mathbf{x}_{j,0}$ is the target transport guidance.
@@ 3.6 Implementation Details
Figure 5: Base Clothing (BC) Standardization. We apply different strategies based on the data. For DeepFashion [deepfashion] and Commerce images [Zhu2024MMVTO, Cheng2023TryOnDiffusion], we generate the base garment while preserving pose and identity via pixel-aligned editing (Prompt V1). For FFHQ [ffhq], we use generative outpainting to expand limited face crops into full-body samples (Prompt V2).
Base Clothing (BC) Dataset. We construct a standardized dataset by processing paired DeepFashion [deepfashion] data, unpaired FFHQ [ffhq] faces, and commercial images [Zhu2024MMVTO, Cheng2023TryOnDiffusion] using Gemini 2.5 Flash Image [nanobanana]. As shown in Figure 5, we employ distinct strategies: pixel-aligned editing for full-body inputs to standardize garments to a black tank top and shorts, and generative outpainting to expand face crops into full-body portraits. The prompts used for this standardization are provided in the Appendix.
Training Setup. Our dataset totals approximately 500k samples: 470K single unpaired images [ffhq, Zhu2024MMVTO, Cheng2023TryOnDiffusion], and 30K paired images from the DeepFashion dataset [deepfashion] (We restrict training to a subset of DeepFashion data to prevent overfitting to its limited number of unique identities ($\approx 100$), while ensuring all identities are still represented.)
Optimization. We use Flux.1 [dev]  [flux2024] as our backbone, training rank-128 LoRA adapters with AdamW with a learning rate of $10^{-4}$. Training runs for 100K iterations on 128 TPUv5 chips with a batch size of 128.
@@ 4 Experiments
4.1 Datasets and Metrics for Evaluation
Following prior work [MCLD, CFLD, Zhou2025LEFFA], we use the DeepFashion In-Shop Clothes Retrieval test split [deepfashion] (8570 image pairs). To assess generalization to in-the-wild scenarios, we additionally evaluate on WPose dataset [Li_UniHuman_2024] (2305 image pairs).
Since our method canonicalizes clothing, direct comparison against ground-truth images (which include diverse garments and backgrounds) is not meaningful. For image similarity (PSNR, SSIM [ssim-iqa]) and perceptual metrics (FID [FID], LPIPS [lpips]), we therefore compare our outputs with the BC version. We further report these metrics on the face regions of the original image in the appendix.
Pose, Identity, Semantic and Perceptual Metrics. To measure how well the generated avatar matches the target pose, we compute Object Keypoint Similarity (OKS) [lin2014microsoft] between the predicted and ground-truth keypoints. Identity fidelity is evaluated within the facial region. We compute Face Similarity (FaceSim) using cosine similarity of ArcFace [arcface] identity embeddings. We report DINOv2 similarity [oquab2023dinov2] as a measure of semantic and structural alignment. For reference-free perceptual quality, we report HPSv3 [Ma_2025_ICCV].
DeepFashion (In-Domain)
[…]
CFLD [CFLD] 17.65 7.15 0.748 0.182 0.48 0.3180 0.9731 4.15
MCLD [MCLD] 18.21 7.08 0.756 0.176 0.49 0.3440 0.9654 4.29
LEFFA [Zhou2025LEFFA] 14.02 4.23 0.755 0.119 0.44 0.5794 0.9409 4.41
OnePoseTrans [oneposetrans] 13.57 8.74 0.605 0.307 0.46 0.5750 0.9476 4.32
UniHuman [Li_UniHuman_2024] 14.05 6.25 0.796 0.156 0.46 0.5810 0.9434 4.03
Gemini 2.5 Flash Image [nanobanana] 16.98 4.59 0.738 0.179 0.43 0.5815 0.9691 7.19
Unpaired Only 15.66 6.54 0.715 0.201 0.47 0.3585 0.9259 4.25
Paired Only 19.38 4.19 0.815 0.071 0.48 0.6255 0.9761 7.25
Ours (Unpaired + paired) 19.36 4.24 0.818 0.075 0.48 0.6047 0.9759 7.24
WPose (Out-of-Domain)
[…]
CFLD [CFLD] 15.43 96.07 0.744 0.208 0.31 0.0885 0.6412 1.94
MCLD [MCLD] 15.64 94.23 0.759 0.201 0.35 0.0995 0.6478 1.96
LEFFA [Zhou2025LEFFA] 16.71 67.85 0.776 0.193 0.32 0.0914 0.5725 2.01
OnePoseTrans [oneposetrans] 17.23 27.43 0.818 0.151 0.33 0.1735 0.7205 4.44
UniHuman [Li_UniHuman_2024] 17.64 27.75 0.807 0.161 0.34 0.1121 0.7207 2.89
Gemini 2.5 Flash Image [nanobanana] 16.67 9.55 0.779 0.149 0.32 0.4713 0.7005 7.35
Unpaired Only 16.13 6.95 0.761 0.215 0.37 0.3805 0.6779 4.37
Paired Only 18.30 6.65 0.820 0.155 0.34 0.4959 0.6972 7.30
Ours (Unpaired + paired) 19.95 5.99 0.860 0.121 0.38 0.5571 0.7394 7.55
Table 1: Quantitative evaluation and ablation study. We report results on DeepFashion (top) and WPose (bottom). The upper section of each block compares our method against state-of-the-art baselines, while the lower section details our ablation study (Unpaired Only, Paired Only, and our full hybrid model). All Pro-Pose results are zero-shot (without test-time fine-tuning). Red indicates the best performance and Blue indicates the second best across all methods.
Figure 6: Identity Preservation vs. Pose Difficulty. FaceSim scores stratified by pose-change magnitude. Pro-Pose closely tracks Ground Truth identity variance across all difficulty bins, while baselines degrade rapidly. With identity breakage defined as FaceSim <0.4<0.4, Pro-Pose fails in only $13.5\%$ of cases vs. $97\%$ for all baselines.
4.2 Quantitative and Qualitative Comparison
We compare our method with recent diffusion-based pose-conditioned avatar generation models, including MCLD [MCLD], CFLD [CFLD], LEFFA [Zhou2025LEFFA], OnePoseTrans [oneposetrans], UniHuman [Li_UniHuman_2024], and Gemini 2.5 Flash Image [nanobanana].
As detailed in Table 1, our method achieves SOTA performance on both benchmarks. On the in-domain DeepFashion set [deepfashion], our model sets a new SOTA, significantly outperforming all baselines across image fidelity (PSNR, SSIM, LPIPS), identity (FaceSim), and perceptual quality (HPSv3). For the in-the-wild WPose dataset [Li_UniHuman_2024], where we use foreground-masked metrics (M-PSNR, M-SSIM, M-LPIPS), our full model (Unpaired + Paired) conclusively dominates all SOTA baselines. This highlights the critical role of our unpaired data strategy in achieving robustness for challenging, in-the-wild data.
Figure 12 provides a qualitative comparison with SOTA methods. As demonstrated, our method generates high-fidelity avatars that accurately preserve the input person’s identity, facial features, and body characteristics. For instance, the eighth row shows robust preservation of face and beard shape. Furthermore, our approach more robustly follows the target pose than any other method (e.g., sixth row) and maintains body shape fidelity relative to the original image (e.g., ninth row).
Stratified Identity Analysis. To quantify identity robustness under varying difficulty, we stratify FaceSim scores by the magnitude of pose change between source and target. As shown in Figure 6, Pro-Pose closely tracks the inherent identity variance of Ground Truth images across all pose-change bins, whereas all baselines exhibit rapid degradation as pose difficulty increases. Defining identity breakage as a FaceSim score below $0.4$, Pro-Pose exhibits a failure rate of only $13.5\%$, while all competing methods (LEFFA, MCLD, UniHuman, OnePoseTrans) fail to preserve identity in over $97\%$ of generated samples.
4.3 Ablation Study
Figure 8: Finetuning Qualitative results. Although the input image suffered from motion blur, leading to blurry base model results, finetuning successfully restored crisp face features, demonstrating the model’s ability to learn robustly from multiple images rather than relying solely on one input.
@@ 4.3 Ablation Study
Ours (Zero-Shot) 18.59 6.88 0.823 0.4837 0.689 7.15
Ours (Fine-Tuned) 19.34 6.76 0.835 0.5722 0.705 7.17
Table 2: Ablation study (Fine-Tuning). Test-time fine-tuning with a few reference images improves identity preservation over our zero-shot base model across all metrics.
@@
As shown in Table 1, while the Paired Only model performs well on in-domain data (DeepFashion [deepfashion]), it suffers on the out-of-domain WPose set [Li_UniHuman_2024], confirming its poor generalization. Figure 7 demonstrates that training solely on the limited paired dataset leads to overfitting and identity leaking from the training data. Alternatively, the model trained exclusively on unpaired data fails when subjected to significant pose changes. Our full method (Unpaired + Paired) synergizes these sources, effectively utilizing the larger dataset to learn robust identity priors while maintaining stable pose guidance.
@@
Figure 9: Emergent Capabilities. $T_{p}$: Edited texture map. By introducing a localized modification — in this case, adding a heart tattoo to the base texture — the edit geometrically propagates with high fidelity to novel target poses $\tilde{p}$ as shown in $I^{Edited}_{\tilde{p}}$. $I_{\tilde{p}}$ shows the output using the un-edited texture map. This demonstrates our model’s capability to apply texture changes such as makeup or tattoos in synthesizing the output portraits.
4.4 Downstream Application: Virtual Try-On
A key advantage of our canonical representation is its utility as a preprocessing stage for downstream tasks. We demonstrate this by combining Pro-Pose with the state-of-the-art Google Virtual Try-On tool [googlevto2023]. As shown in Figure 10, applying VTO directly to in-the-wild images can lead to erroneous results due to challenging poses, cluttered backgrounds, and complex clothing. In contrast, first generating a clean, reposed canonical avatar with Pro-Pose significantly improves VTO fidelity.
Figure 10: Downstream Application: Virtual Try-On. Applying VTO [googlevto2023] directly to an unconstrained in-the-wild image (VTO on RI) yields limited fidelity. Using Pro-Pose to first generate a clean, reposed canonical avatar (Pro-Posed RI) significantly improves VTO quality (VTO on PRI).
4.5 Emergent Model Properties: Direct Texture Editing.
Our framework unlocks other powerful features without explicit training such as direct texture editing. Because Pro-Pose extracts a canonical UV map, any manual edits to this texture (e.g., adding a tattoo) propagate with strict geometric consistency to the final reposed avatar, preserving identity without distortion (Figure 9 adding a heart tattoo on texture map can transfer to the reposed person).
Figure 11: Limitations of the base model. Left: Extreme pose changes (e.g., back-to-front) result in minimal UV texture overlap, forcing the model to hallucinate unseen regions. While general attributes like beard shape or hair color may be inferred from partial cues, precise facial structure can lose fidelity. Right: Zero-shot synthesis is constrained by the accuracy of single-view 3D body estimation. Errors in SMPL-X fitting can propagate during the reposing stage, leading to inaccurate body proportions (e.g., a noticeable reduction in thigh volume in the ProPose results compared to the ground truth).
@@ 5 Limitations and Future Work
While our framework significantly advances single-image avatar reposing, some limitations remain. First, the base model struggles with extreme pose changes, such as synthesizing a frontal pose from a single rear-view image. Consequently, the model inevitably hallucinates facial features (Figure 11, left). However, this limitation can be mitigated by providing multiple reference images of the subject and applying our fine-tuning approach. Second, our method relies heavily on the accuracy of the underlying SMPL-X body estimation. During zero-shot in-the-wild evaluation, we extract body shape parameters exclusively from the single reference image and combine them with the pose parameters of the target. However, recovering 3D body shape from a single image can be erroneous. Figure 11 (right) showcases an example from the WPose dataset where the reference body shape is inaccurately estimated. When this flawed shape geometry is reposed into a novel target view, the errors propagate directly to the generated canonical avatar, resulting in incorrect body shape. This highlights that our synthesis fidelity is ultimately bounded by the robustness of single-view 3D body priors. Third, while our full-body UV conditioning preserves visible skin details, self-occluded body regions cannot be faithfully reconstructed from a single view. Future work will explore more robust body estimation techniques and extend the framework to zero-shot multi-reference consistency.
@@ 6 Conclusion
In this paper, we presented a novel framework for generating photorealistic, reposed human avatars from unconstrained single images. By introducing a self-supervised Donor-based UV Reposing strategy operating on full-body UV textures, we effectively decoupled pose from appearance, enabling our Flow Matching generator to learn robust full-body identity preservation from massive unpaired image collections. We further showed that this strong base model can be rapidly specialized via multi-image fine-tuning to handle more extreme reposing scenarios. By standardizing data into a base clothing and formulating generation in canonical UV space, our approach provides a scalable foundation for creating high-fidelity, controllable digital avatars from everyday photography, and serves as a powerful canonical intermediate for downstream applications such as virtual try-on.
Figure 12: Qualitative Comparison. Results are shown for the DeepFashion [deepfashion] (rows 1-3) and WPose [Li_UniHuman_2024] (bottom rows) datasets. The columns show the Original Person, Target Pose, and results from state-of-the-art methods (MCLD [MCLD], Leffa [Zhou2025LEFFA], UniHuman [Li_UniHuman_2024], OnePoseTrans [oneposetrans], Gemini 2.5 Flash Image [nanobanana]) compared to Ours (Zero-Shot) and the Ground Truth (GT). Our method demonstrates superior performance in preserving the person’s identity, particularly the facial characteristics, across varying poses. For instance, in the third and fifth rows, previous methods introduce significant identity distortion and loss of facial likeness (e.g., changes in jawline or features). In contrast, Ours robustly preserves the unique face structure and identity of the original person, closely matching the GT result. (prompt used for Gemini 2.5 Flash Image: A professional studio portrait of the person from IMAGE_1, maintaining their exact facial features, hairstyle, build, and identity. They are captured in the pose and camera angle shown in IMAGE_2).
@@
Appendix 0.A Base Clothing (BC) Dataset.
To train our model, we utilize paired data from DeepFashion [deepfashion], unpaired single images from FFHQ [ffhq], and commercial data derived from e-commerce websites [Zhu2024MMVTO, Cheng2023TryOnDiffusion]. To ensure visual uniformity across these diverse sources, we generate the Base Clothing (BC) dataset using an identity-preserving I2I model (Gemini 2.5 Flash Image [nanobanana]), as shown in Figure 5 (main paper). Depending on the nature of the input data, we employ either Prompt V1 or Prompt V2:
@@
Precisely edit IMAGE_0 to change the garment worn by the person in the photo to a black sleeveless tank top and 3-inch black shorts. Do not change the person in the image or the framing. Remove or replace any accessories such as bags, hats, jewelry and sunglasses in IMAGE_0. Do not alter IMAGE_0 in any other way. Keep it pixel-aligned.
@@
Appendix 0.B Canonical UV-Space Formulation